OpenStudy (anonymous):

Ismael is comparing cell phone plans before upgrading his phone. Ameri-Mobile offers a low activation fee, but a high monthly payment. Cell-U-Later offers a lower monthly rate, but the activation fee is higher. Create a possible algebraic expression for both Ameri-Mobile and Cell-U-Later that shows the amount paid after an unknown amount of months have passed. Justify how you created those expressions, and identify what each term and factor represents in terms of the cell phone plans

Mehek (mehek14):

is this Algebra 1?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

please help

OpenStudy (lilshane):

AM-20+100x am=ameri-mobile (af=activation fee) (MP=monthly payment) CUL+300-50x (CUL=Cell-U-Later)(xM=unknown months)i created the variables by using the initials of the words thats a equation that i used on the 01.04 quiz

OpenStudy (anonymous):

thankssss

OpenStudy (lilshane):

i hoped tht helped if so can i get a medal

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know how you made those number up? does it matter which number you use? i don't wanna get a bad grade is all

OpenStudy (anonymous):

and what property do you use